# Disturbance-Robust Backup Control Barrier Functions: Safety Under Uncertain Dynamics

### v1

@@ Abstract
Obtaining a controlled invariant set is crucial for safety-critical control with control barrier functions (CBFs) but is non-trivial for complex nonlinear systems and constraints. Backup control barrier functions allow such sets to be constructed online in a computationally tractable manner by examining the evolution (or flow) of the system under a known backup control law. However, for systems with unmodeled disturbances, this flow cannot be directly computed, making the current methods inadequate for assuring safety in these scenarios. To address this gap, we leverage bounds on the nominal and disturbed flow to compute a forward invariant set online by ensuring safety of an expanding norm ball tube centered around the nominal system evolution. We prove that this set results in robust control constraints which guarantee safety of the disturbed system via our Disturbance-Robust Backup Control Barrier Function (DR-BCBF) solution. Additionally, the efficacy of the proposed framework is demonstrated in simulation, applied to a double integrator problem and a rigid body spacecraft rotation problem with rate constraints.
@@ I Introduction
Control barrier functions (CBFs) [1, 2], are a popular approach to assuring safety of autonomous systems by encoding notions of safety directly into existing controllers and providing sufficient conditions for forward invariance of safe sets. However, obtaining safe sets for which every point has a safe control action (i.e., is controlled invariant) is difficult for high-dimensional systems. Additionally, dynamics models are seldom perfect. In this paper we seek to solve both of these problems simultaneously.
@@
To address the problem of controlled invariance, we adapt the backup set method from [3, 4] based on online backward reachability. This method establishes a controlled invariant safe set implicitly using the flow of the system under a prescribed backup control law. This approach is computationally tractable even for complex systems. For affine nonlinear systems, this technique generates linear control constraints which can be used to efficiently solve for point-wise optimal control signals for an arbitrary primary controller.
The second problem we address is that of model uncertainty, in the form of unknown bounded disturbances. Much work has been done in recent years in the CBF literature to address model uncertainty and disturbances. In [5] robust CBFs (RCBFs) are introduced which account for the worst-case disturbance. RCBFs with input constraints were considered in [6] with an application to spacecraft control. In [7] the authors synthesized RCBFs and control Lyapunov functions together and provided additional conditions for convergence guarantees. RCBFs for relative degree two, sampled-data systems were developed and applied to robotic grasping in [8]. However, these approaches all assume that an explicit representation of a controlled invariant safe set can be found. Unlike these approaches, [9] leverages the mixed-monotonicity property of certain systems to over-approximate the reachable sets computed using the backup set method.
@@
Figure 1: Overview of the proposed disturbance-robust safety-critical control framework. $\mathcal{C}_{\rm I}$ represents a forward invariant subset of an unknown controlled invariant set $\mathcal{C}_{\rm D}$ and guarantees safety of the disturbed system.
This manuscript presents a robust safety-critical control framework, displayed in Figure 1, which extends the method of backup CBFs to make formal safety guarantees on systems with unknown disturbances. The main contributions of this paper are two-fold. First, we introduce disturbance-robust backup CBFs from which we derive tractable robust safety conditions which can be synthesized directly into existing controllers. Second, we demonstrate the effectiveness of the approach using two numerical simulations: a toy problem and a spacecraft rotation example.
@@ II Preliminaries
This section reviews the fundamental concepts of safety-critical control using control barrier functions (CBFs) and constructing invariant sets using an implicit approach, also referred to as the backup set method.
@@ II-A Control Barrier Functions
Consider a nonlinear control affine dynamical system modeled as
$\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u},$ (1)
where $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$ and $g:\mathcal{X}\rightarrow\mathbb{R}^{n\times m}$ are Lipschitz continuous functions, $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ represents the state vector, and $\boldsymbol{u}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ represents the control vector. $\mathcal{X}$ is the set of all possible states and $\mathcal{U}$ is the admissible control set. For an initial condition $\boldsymbol{x}(0)=\boldsymbol{x}_{0}\in\mathcal{X}$ if $\boldsymbol{u}$ is given by a locally Lipschitz feedback controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$, the closed-loop system (1) has a unique solution $\boldsymbol{\phi}^{n}(t,\boldsymbol{x}_{0})$ over an interval of existence.
In the context of this work, safety is defined as a region of the state space, known as the safe set $\mathcal{C}_{\rm S}$, within which the dynamical system must stay. Safe controllers are ones that render this safe set forward invariant. A set $\mathcal{C}\subset\mathbb{R}^{n}$ is forward invariant along (1) if $\boldsymbol{x}_{0}\in\mathcal{C}\implies\boldsymbol{\phi}^{n}(t,\boldsymbol{x}_{0})\in\mathcal{C},$ for all $t>0$. Now, consider the safe set $\mathcal{C}_{\rm S}$ as the 0-superlevel set of a continuously differentiable function $h:\mathcal{X}\rightarrow\mathbb{R}$ with
$\mathcal{C}_{\rm S}\triangleq\{\boldsymbol{x}\in\mathcal{X}:h(\boldsymbol{x})\geq 0\},$ (2)
and the gradient along the boundary of $\mathcal{C}_{\rm S}$ remains nonzero, i.e., $h(\boldsymbol{x})=0\implies\nabla h(\boldsymbol{x})\neq 0$.
Definition 1.
@@
where $L_{f}h$ and $L_{g}h$ are the Lie derivatives of $h$ along $f$ and $g$, respectively.
Finally, we arrive at the main result of [1], relating CBFs and forward invariance to safety.
Theorem 1.
@@
for all $\boldsymbol{x}\in\mathcal{C}_{\rm S}$ will render the set $\mathcal{C}_{\rm S}$ forward invariant.
The sufficient condition for forward invariance of the safe set is linear in the control, often motivating the use of CBFs to supply safety constraints to a quadratic program (QP). For an arbitrary primary controller $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, it is possible to ensure safety of (1) by solving the following point-wise optimization problem for the safe control, $\boldsymbol{u}_{\rm safe}$:
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad\quad(\text{CBF-QP})\\ \quad\quad\quad\quad\text{s.t.}\quad L_{f}h(\boldsymbol{x})+L_{g}h(\boldsymbol{x})\boldsymbol{u}\geq-\alpha(h(\boldsymbol{x})).\end{gathered}$
A key challenge is actually obtaining an explicit representation of such a function $h$ where a safe control signal satisfying (3) can always be found. Depending on the safe set this may be difficult or impossible, especially for high dimensional systems. This therefore motivates the use of an extension of CBFs known as backup CBFs.
II-B Implicitly Defined Invariant Sets
First introduced in [3] and expanded upon in [4], the backup set approach relies on filtering with respect to an implicitly defined controlled invariant set.
Definition 2.
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is controlled invariant if there exists a controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$ which renders $\mathcal{C}$ forward invariant for a dynamical system defined by (1).
To construct an implicit controlled invariant set, first assume that we have defined a set $\mathcal{C}_{\rm S}$ describing our state constraints, which is not necessarily controlled invariant. Now suppose that there exists a set within $\mathcal{C}_{\rm S}$ which we call a backup set, $\mathcal{C}_{\rm B}$, such that $\mathcal{C}_{\rm B}\subset\mathcal{C}_{\rm S}$. This set is defined similar to (2) with a continuously differentiable function $h_{\rm b}$, it is known to be controlled invariant, and it is made forward invariant by a backup control law defined by $\boldsymbol{u}_{\rm b}:\mathcal{X}\rightarrow\mathcal{U}$. This backup control law and set can often be obtained via an application of Lyapunov stability theory around local regions of attraction [10]. The closed-loop system under such a backup control law is denoted
@@ II-B Implicitly Defined Invariant Sets
To obtain this implicitly defined controlled invariant set, $\mathcal{C}_{\rm BI}$, satisfying $\mathcal{C}_{\rm B}\subseteq\mathcal{C}_{\rm BI}\subseteq\mathcal{C}_{\rm S}$, one must ensure that the trajectory of the system under $\boldsymbol{u}_{\rm b}(\boldsymbol{x})$ remains in $\mathcal{C}_{\rm S}$ over a finite time horizon, and that the final point in the trajectory lies within the backup set $\mathcal{C}_{\rm B}$. It is assumed that for any $\boldsymbol{x}\in\mathcal{X}$ there exists a unique solution $\boldsymbol{\phi}^{n}_{\rm b}:[0,T]\times\mathcal{X}\rightarrow\mathcal{X}$ which satisfies:
@@
The solution is the flow of the system over the interval $[0,T]$ for $T\in\mathbb{R}_{>0}$ starting at state $\boldsymbol{x}$ under the backup control law $\boldsymbol{u}_{\rm b}$. Therefore $\mathcal{C}_{\rm BI}$ is defined as
@@
While $\mathcal{C}_{\rm BI}$ is seldom the largest controlled invariant subset of $\mathcal{C}_{\rm S}$ (known as the viability kernel [11]), it is an under-approximation of such a set which can be computed in real-time.
As before, these nonlinear inequalities on the state which define the controlled invariant set $\mathcal{C}_{\rm BI}$ can be converted into linear control constraints. We require $\forall\mkern 2.0mu\tau\in[0,T]$, and class-$\mathcal{K}_{\infty}$ functions $\alpha$ and $\alpha_{\rm b}$ that
$\dot{h}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}),\boldsymbol{u})$ $\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))),$ (9a)
$\dot{h}_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}),\boldsymbol{u})$ $\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))).$ (9b)
Expanding the total derivative terms, the constraints can be written as
$\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\frac{\partial\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})}{\partial\boldsymbol{x}}\dot{\boldsymbol{x}}$ $\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))),$ (10a)
$\!\!\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\frac{\partial\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x})}{\partial\boldsymbol{x}}\dot{\boldsymbol{x}}$ $\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))),$ (10b)
for all ${\tau\in[0,T]}$, where ${\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}}$. Here, ${\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\triangleq\partial\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})/\partial\boldsymbol{x}}$ is the sensitivity matrix, or state-transition matrix (STM), which captures the sensitivity of the flow to perturbations in the initial condition $\boldsymbol{x}$. The sensitivity matrix is the solution to the differential equation
$\begin{gathered}{\dot{\boldsymbol{\Phi}}}_{\rm b}^{n}(\tau,\boldsymbol{x})=F_{\rm cl}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x}),\\ \mkern 9.0mu\boldsymbol{\Phi}_{\rm b}^{n}(0,\boldsymbol{x})=\boldsymbol{I}_{n\times n},\end{gathered}$ (11)
where $F_{\rm cl}$ is the Jacobian of the closed-loop backup dynamics (4) evaluated at $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and $\boldsymbol{I}_{n\times n}$ is the $n\times n$ identity matrix. This restricts the backup controller to be smooth. While in some special cases a closed-form solution of the STM is known, in general the STM can be numerically integrated along with the backup flow.
The inequality constraints given by (10a) represent an infinite number of constraints. Therefore, to construct a tractable optimization problem they are discretized and enforced at discrete times along the flow. To ensure safety between sample points, (10a) is tightened using a constant $\varepsilon_{\Delta}$:
@@
As before, the safety of (1) can be enforced for an arbitrary primary controller, $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, by solving the backup CBF optimization problem for the safe control with the newly tightened, implicitly defined constraints:
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad(\text{Backup CBF-QP})\\ \text{s.t. }\\ \nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\dot{\boldsymbol{x}}\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\varepsilon_{\Delta}),\\ \mkern-6.0mu\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\dot{\boldsymbol{x}}\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))),\end{gathered}$
for all $\tau\in\{0,\Delta,\dots,T\}$, where ${\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}}$.

### v2

@@ Abstract
Obtaining a controlled invariant set is crucial for safety-critical control with control barrier functions (CBFs) but is non-trivial for complex nonlinear systems and constraints. Backup control barrier functions allow such sets to be constructed online in a computationally tractable manner by examining the evolution (or flow) of the system under a known backup control law. However, for systems with unmodeled disturbances, this flow cannot be directly computed, making the current methods inadequate for assuring safety in these scenarios. To address this gap, we leverage bounds on the nominal and disturbed flow to compute a forward invariant set online by ensuring safety of an expanding norm ball tube centered around the nominal system evolution. We prove that this set results in robust control constraints which guarantee safety of the disturbed system via our Disturbance-Robust Backup Control Barrier Function (DR-bCBF) solution. The efficacy of the proposed framework is demonstrated in simulation, applied to a double integrator problem and a rigid body spacecraft rotation problem with rate constraints.
@@ I Introduction
Control barrier functions (CBFs) [1], are a popular approach to assuring safety of autonomous systems by encoding safety into existing controllers and providing sufficient conditions for forward invariance of safe sets. However, obtaining safe sets for which every state has a safe control action (a.k.a. controlled invariant sets) is difficult for high-dimensional systems, especially when considering input bounds. Additionally, dynamics models are seldom perfect. In this letter we seek to solve both of these problems simultaneously.
@@
To address the problem of controlled invariance, we adapt the backup set method [2, 3, 4] based on online backward reachability. This method establishes a controlled invariant safe set implicitly using the flow of the system under a prescribed backup control law. This approach is computationally tractable even for complex systems. For affine nonlinear systems, this technique generates linear control constraints which can be used to efficiently solve for point-wise optimal control signals for an arbitrary primary controller.
The second problem we address is that of model uncertainty, which has been studied extensively in the CBF literature. Robust methods [5, 6, 7, 8] typically rely on accounting for worst-case disturbances through an upper disturbance bound, and these methods can be made less conservative via disturbance estimation [9]. The notion of input-to-state safety, defined first in [10] and extended for CBFs in [11], provides a technique for handling input disturbances and has been successfully applied in multiple scenarios [12, 13]. Adaptive CBF methods have been shown to assure safety in the presence of parametric dynamics uncertainty [14, 15] and a robust adaptive CBF extension can reduce conservatism and closed-loop chattering [16]. Learning [17, 18] and data-driven [19, 20, 21] approaches have also been developed to account for uncertainty in dynamics, state, or both. Lastly, for mixed-monotone systems their decomposable structure can be exploited to produce robustly forward invariant sets [22]. While these approaches present viable solutions to addressing model uncertainty, they assume that a controlled invariant safe set can be found explicitly—a strong assumption for many systems and safety constraints. Works [23] and [24] do not make this assumption, but the former is specific to mixed-monotone systems, and the latter assumes perfect dynamics knowledge and bounded measurement error.
@@
Figure 1: Depiction of the proposed disturbance-robust safety-critical control framework. $\mathcal{C}_{\rm I}$ represents a forward invariant subset of an unknown controlled invariant set $\mathcal{C}_{\rm D}$ and guarantees safety of the disturbed system.
The main contribution of this work is a novel approach to address controlled invariance and dynamics disturbances simultaneously through the formulation of disturbance-robust backup CBFs. Unlike existing works, a controlled invariant set describing safety is not assumed to be known a priori, but is instead constructed online. We first derive forward invariance conditions for a subset inside a controlled invariant set of the disturbed system (displayed in Figure 1). Then we robustify these conditions and integrate them with existing controllers via a quadratic program. The proposed framework guarantees safety for a broad class of nonlinear systems with limited control authority even in the presence of unknown, bounded disturbances. We demonstrate the effectiveness of the approach using two numerical simulations: an illustrative double integrator system and a spacecraft rotation example.
@@ II-A Control Barrier Functions
Consider a nonlinear control affine system of the form
$\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u},\mkern 8.0mu{\color[rgb]{0,0,0}\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{n},\mkern 8.0mu\boldsymbol{u}\in\mathcal{U}\subseteq\mathbb{R}^{m},}$ (1)
where $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$ and $g:\mathcal{X}\rightarrow\mathbb{R}^{n\times m}$ are Lipschitz continuous functions. It is assumed that $\mathcal{U}$ is an $m$-dimensional convex polytope. For an initial condition $\boldsymbol{x}(0)=\boldsymbol{x}_{0}\in\mathcal{X}$ if $\boldsymbol{u}$ is given by a locally Lipschitz feedback controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$, the closed-loop system (1) has a unique solution $\boldsymbol{\phi}^{n}(t,\boldsymbol{x}_{0})$ over an interval of existence.
@@
where $L_{(\cdot)}h$ is the Lie derivative of $h$ along function ()($\cdot$)( ⋅ ).
Theorem 1 ([1]).
@@
for all $\boldsymbol{x}\in\mathcal{C}_{\rm S}$ renders the set $\mathcal{C}_{\rm S}$ forward invariant.
For an arbitrary primary controller, $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, it is possible to ensure the safety of (1) by solving the following point-wise optimization problem for the safe control, $\boldsymbol{u}_{\rm safe}$:
$\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu$ $\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}$ (CBF-QP)
s.t. $L_{f}h(\boldsymbol{x})+L_{g}h(\boldsymbol{x})\boldsymbol{u}\geq-\alpha(h(\boldsymbol{x})).$
A key challenge is obtaining an explicit representation of such a function $h$ where a safe control signal satisfying (2) can always be found. Depending on the safe set this may be difficult or impossible, especially for high dimensional systems. This therefore motivates the use of an extension of CBFs known as backup CBFs.
II-B Implicitly Defined Controlled Invariant Sets
First introduced in [2] and expanded upon in [3], the backup CBF approach relies on obtaining an implicitly defined controlled invariant set. A set $\mathcal{C}\subset\mathbb{R}^{n}$ is controlled invariant if there exists a controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$ which renders $\mathcal{C}$ forward invariant for (1).
To construct an implicit controlled invariant set, first assume that we have defined a set $\mathcal{C}_{\rm S}$ describing our state constraints, which is not necessarily controlled invariant. Now suppose that there exists a set within $\mathcal{C}_{\rm S}$ which we call a backup set, $\mathcal{C}_{\rm B}$, such that $\mathcal{C}_{\rm B}\subset\mathcal{C}_{\rm S}$. This set is defined similar to $\mathcal{C}_{\rm S}$ with a continuously differentiable function $h_{\rm b}$, it is known to be controlled invariant, and it is made forward invariant by a backup control law defined by $\boldsymbol{u}_{\rm b}:\mathcal{X}\rightarrow\mathcal{U}$. The closed-loop system under $\boldsymbol{u}_{\rm b}$ is denoted as
@@ II-B Implicitly Defined Controlled Invariant Sets
It is assumed that for any $\boldsymbol{x}\in\mathcal{X}$ there exists a unique solution $\boldsymbol{\phi}^{n}_{\rm b}:[0,T]\times\mathcal{X}\rightarrow\mathcal{X}$ which satisfies:
@@
The solution is the flow of the system over the interval $[0,T]$ for $T\in\mathbb{R}_{>0}$ starting at state $\boldsymbol{x}$ under the backup control law $\boldsymbol{u}_{\rm b}$. To obtain an implicitly defined controlled invariant set, $\mathcal{C}_{\rm BI}$, satisfying $\mathcal{C}_{\rm B}\subseteq\mathcal{C}_{\rm BI}\subseteq\mathcal{C}_{\rm S}$, one must ensure that the trajectory of the system under $\boldsymbol{u}_{\rm b}(\boldsymbol{x})$ remains in $\mathcal{C}_{\rm S}$ over a finite time horizon, and that the final point in the trajectory lies within the backup set $\mathcal{C}_{\rm B}$. Therefore $\mathcal{C}_{\rm BI}$ is defined as
@@
The sufficient condition for forward invariance of $\mathcal{C}_{\rm BI}$, and thus safety with respect to $\mathcal{C}_{\rm S}$, is then
$\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\dot{\boldsymbol{x}}$ $\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))),$ (8a)
$\!\!\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\dot{\boldsymbol{x}}$ $\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))),$ (8b)
for all ${\tau\in[0,T]}$ and class-$\mathcal{K}_{\infty}$ functions $\alpha$ and $\alpha_{\rm b}$. Here, ${\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}}$ and ${\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\triangleq\partial\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})/\partial\boldsymbol{x}}$ is the sensitivity matrix, or state-transition matrix (STM), which captures the sensitivity of the flow to perturbations in the initial condition $\boldsymbol{x}$. The STM is the solution to
$\begin{gathered}{\dot{\boldsymbol{\Phi}}}_{\rm b}^{n}(\tau,\boldsymbol{x})=F_{\rm cl}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x}),\mkern 9.0mu\boldsymbol{\Phi}_{\rm b}^{n}(0,\boldsymbol{x})={\color[rgb]{0,0,0}\boldsymbol{I}},\end{gathered}$ (9)
where $F_{\rm cl}$ is the Jacobian of the closed-loop backup dynamics (3) evaluated at $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and ${\color[rgb]{0,0,0}\boldsymbol{I}}$ is the $n\times n$ identity matrix.
Because the inequality in (8a) represents an infinite number of constraints, in practice these are discretized and enforced at discrete times along the flow. To ensure safety between sample points, (8a) is tightened via a constant $\varepsilon_{\Delta}$[4, Thm. 3]:
@@
As in (CBF-QP), the safety of (1) can be enforced for a primary controller, $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, by solving an optimization problem for the safe control with constraints (8), where the right-hand side of (8a) is replaced by $-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\varepsilon_{\Delta})$.